2.) Estimate the population proportions for each of the following data:


a.) In a plant box consisting of 120 seedlings, 80 seedlings were treated with growth enhancer. Estimate p and q.

b.) In a survey of 80 children, 48 like to watch horror films. Find and , where is the proportion of children who like to watch horror films.


1
Expert's answer
2021-06-07T11:57:42-0400

2.

a.


"p=\dfrac{x}{n}=\dfrac{80}{120}=\dfrac{2}{3}""q=1-p=1-\dfrac{2}{3}=\dfrac{1}{3}"



b.


"p=\dfrac{x}{n}=\dfrac{48}{80}=0.6""q=1-p=1-0.6=0.4"